OpenStudy (candaycorn):

Someone please help! Suppose you have 18 more quarters than dimes. You have a total of $8.00. How many of each coin do you have? Use the subsitution method..

OpenStudy (opti):

Make a system of equations

OpenStudy (mathstudent55):

Use one equation to deal with how many coins of each type you have. Use the other equation to deal with how much money you have.

OpenStudy (imstuck):

When you make a system of equations for this, base either the number of quarters you have on the number of dimes or the other way around. If you base it on the number of dimes, you have that the number of quarters is dimes plus 18.

OpenStudy (imstuck):

That would then be your first equation. Up above, mathstudent55 told you that one equation was based on the number of coins while the other was based on what they are worth; one is money value one is number of coins. This one we just did was number of coins.

OpenStudy (imstuck):

So your equation is q = d + 18 Do you understand that so far?

OpenStudy (imstuck):

Now we have another part of that problem that deals with what they are worth. We know that a quarter is worth 25 cents and dime is worth 10 cents, and we also know that, added together, the value of the coins we have is $8. We don't know how many we have, but we do know what their individual values are. So the second equation will be based on the fact that a quarter is 25 cents and a dime is 10 cents and added together they equal $8. .25q + .10d = 8.00

OpenStudy (imstuck):

Now you have a system to solve: q = d + 18 .25q + .10d = 8.00

OpenStudy (candaycorn):

Thank you very much, You made that a lot easier, I apreciate your time!:)